OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? 54 27 81 62

OpenStudy (anonymous):

b + g = 89 --> g = 89 - b b = 2g + 8 Now just sub 89 - b in for g and then solve for b and you will have your answer